If I have $4000 and I invested it into two accounts one receives 4% and the other account receives 3 % how much did I put in each account to make $150 annunally?

Answer by ankor@dixie-net.com(22740)

If I have $4000 and I invested it into two accounts one receives 4% and the other account receives 3 % how much did I put in each account to make $150 annunally?
:
let x = amt invested at 4%
the total amt is $4000, therefore
(4000-x) = amt invested at 3%
:
Using the decimal equiv
.04x + .03(4000-x) = 150
.04x + 120 - .03x = 150
.04x - .03x = 150 -120
.01x = 30
x = 30/.01
x = $3000 invested at 4%
then, obviously
4000 - 3000 = $1000 invested at 3%
:
:
Check, find the amt earned by each amt
.04 * 3000 = 120
.03 * 1000 = 30
-----------------
total amt: 150